water will overflow from the bow Constants1 Periodic Table the water level in the bowl will decrease An aluminum bowl with a diameter of 23 em and a height of 6.0 em is completely filled with water. The iritial temperature of the bowl and water is 13 C. The bowl is now placed on a stove burner and heated to 88 C Previous Answers Correct Part B Calculate the volume of water that overflows or the drop in water level in the pan, whichever is appropriate Express your answer using two significant figures Submit Previous Answers Reavest Anewer incorrect; Try Again
Answer #1

a)

Water will overflow from the bowl

b)

V_{o}=\pi r^{2}h=\pi \times \left ( \frac{0.23}{2} \right )^{2}(0.06)=2.5\times 10^{-3}m^{2}

The Volume of water that overflows in the pan is

V_{spill}=\Delta V_{water}-\Delta V_{al}=\left ( \beta _{w}-\beta _{Al} \right )V_{o}\Delta T

V_{spill}=(21-7.2)\times 10^{-5}\times (2.5\times 10^{-3})\times (88-13)=2.58\times 10^{-5}m^{3}